2x+ 4y=16 x+3=5


solution:(14,-3)
need a break down of how my text book got these solutions

Answer by stanbon(75887)

2x+ 4y=16
x+3y =5
------------
Divide thru the 1st equation by 2 to get:
x + 2y = 8
x + 3y = 5
---------------
Subtract the 1st from the 2nd to solve for "y":
y = -3
--------------
Substitute into x + 3y = 5 to solve for "x":
x + 3*-3 = 5
x - 9 = 5
x = 14
---------------
solution:(14,-3)
======================
Cheers,
Stan H.